### v1

# DanceChat: Large Language Model-Guided Music-to-Dance Generation

Abstract.
Music-to-dance generation aims to synthesize human dance motion conditioned on musical input. Despite recent progress, significant challenges remain due to the semantic gap between music and dance motion, as music offers only abstract cues, such as melody, groove, and emotion, without explicitly specifying the physical movements. Moreover, a single piece of music can produce multiple plausible dance interpretations. This one-to-many mapping demands additional guidance, as music alone provides limited information for generating diverse dance movements. The challenge is further amplified by the scarcity of paired music and dance data, which restricts the model’s ability to learn diverse dance patterns. In this paper, we introduce DanceChat, a Large Language Model (LLM)-guided music-to-dance generation approach. We use an LLM as a choreographer that provides textual motion instructions, offering explicit, high-level guidance for dance generation. This approach goes beyond implicit learning from music alone, enabling the model to generate dance that is both more diverse and better aligned with musical styles. Our approach consists of three components: (1) an LLM-based pseudo instruction generation module that produces textual dance guidance based on music style and structure, (2) a multi-modal feature extraction and fusion module that integrates music, rhythm, and textual guidance into a shared representation, and (3) a diffusion-based motion synthesis module together with a multi-modal alignment loss, which ensures that the generated dance is aligned with both musical and textual cues. Extensive experiments on AIST++ and human evaluations show that DanceChat outperforms state-of-the-art methods both qualitatively and quantitatively. The additional qualitative samples can be found at https://dancechat.github.io/anon/.
3D Motion Synthesis, Cross-Modality Learning, Tempo Synchronization
ccs: Computing methodologies Motion processingccs: Applied computing Sound and music computing
Figure 1. Our approach inspired by real-world dance learning. Just as dance learners rely on choreographers to interpret music, DanceChat uses textual dance instructions to bridge the semantic gap between music and motion. Large Language Models (LLMs) act as pseudo-choreographers, transforming musical cues into detailed, actionable guidance for dance generation.
1. Introduction
Dance has long served as a powerful medium of expression in cultures, playing a central role in rituals, social connections, and artistic activities (Carroll and Moore, 2008). In the digital age, dance flourishes on social media, allowing people to share creativity and connect globally. Recent advances in deep learning have enabled computational exploration of this art form, particularly through music-to-dance generation, where models automatically create dance movements synchronized to music. This technology has impactful applications in entertainment, enabling realistic, rhythmically aligned virtual idols and 3D characters in virtual reality. Beyond entertainment, automated dance generation is valuable in fields like robotics, digital human modeling, and physical rehabilitation, where it helps create lifelike movement simulations and supports human-computer interaction.
Most existing music-to-dance generation approaches fall into two main categories: regression models (Huang et al., 2021; Li et al., 2021), which are typically designed to directly learn a mapping from input music to dance motions, and generative models (Siyao et al., 2022; Tseng et al., 2023; Zhang et al., 2024b; Huang et al., 2024), which aim to learn a distribution over plausible dances conditioned solely on the input music. However, both approaches lack explicit movement instructions, and rely only on abstract nature of musical elements such as rhythm, tempo, and mood. This creates a broad semantic gap between music and motion, as illustrated in Fig. 1, and complicates the task of generating physically plausible movements that align with the music’s genre and maintain aesthetic quality. Moreover, the scarcity of labeled dance data, due in part to the high cost of motion capture, makes it difficult for models to learn a well-generalized and diverse motion distribution. With limited training samples, the model struggles to capture the full variability of human dance, resulting in outputs that may lack stylistic diversity or fail to generalize to unseen music.
These challenges highlight the need for advanced methods capable of bridging complex semantic gaps and injecting semantic-rich textual instructions without relying on additional annotations. Recent advancements in large language models (LLMs) have demonstrated remarkable success in capturing nuanced semantic relationships across various domains (Wu et al., 2023; Han et al., 2023; Doh et al., 2024; Gardner et al., 2024; Ding et al., 2024). They are trained on diverse content, including detailed discussions on music, enabling them to interpret musical elements such as tempo, key, and chord progressions. This allows them to understand and generate contextually appropriate descriptions that capture the structure, rhythm, and mood of a piece. This inspires new approaches for tasks involving multi-modal relationships and encourages us to investigate the possibilities for leveraging LLMs to enhance the controllability of music-to-dance generation.
Building on these insights, our method adapts the capabilities of LLMs to address the semantic gap in music-driven dance generation. We propose a novel approach, named DanceChat, which conceptualizes the generation model as a dancing learner and leverages LLMs as a pseudo-choreographer. Specifically, we extract structured musical descriptors, including key, tempo, and chord progressions alongside the dance genre context, and feed this information into LLM. The model then generates corresponding textual dance instructions, which serve as a pseudo-guide to condition the motion generation model during training. This helps reduce the semantic gap by offering interpretative cues that map abstract musical elements to concrete motion patterns.
To further enhance alignment between modalities, we encode the input music into finer-grained descriptors such as beat and downbeat positions, which are known to correlate strongly with movement timing in dance. These rhythmic cues, together with the LLM-generated textual instructions, are integrated through a multi-modal fusion module, enabling the model to form a richer and more coherent understanding of the music. By combining musical structure, rhythmic signals, and semantic instructions, DanceChat produces dance motions that are rhythmically aligned, physically plausible, and stylistically expressive.
In summary, our contributions are three-fold:
1) A novel LLM-guided framework that generates context-driven dance movement instructions aligned with specific musical structure and style.
2) A multi-modal fusion paradigm that integrates beat-aware musical information with the textual dance instructions, enriching the model’s capacity for music-to-dance alignment.
3) A multi-modal alignment loss that leverages text as an intermediary to enhance the semantic connection between music and dance, reducing the gap between these modalities and improving the coherence and expressiveness of generated dance movements.
2. Related Work
Figure 2. Overview of our approach. DanceChat consists of three main components: (a) Pseudo Instruction Generation module extracts music caption (tempo, key, chord, etc.) from the given music condition, (b) Multi-modal Fusion module integrates the encoded representations of music, beats, and text into a unified conditional embedding, and (c) Diffusion-based Motion Synthesis module uses this unified embedding to guide the generation of realistic dance movement.
2.1. 3D Human Motion Generation
Common approaches to 3D human motion generation (Zhu et al., 2023) can be broadly categorized into three sub-tasks: text-conditioned motion generation (Ahn et al., 2018; Tevet et al., 2022a; Chen et al., 2023; Zhang et al., 2022; Kim et al., 2023; Lin et al., 2023; Zhang et al., 2023; Zhou and Wang, 2023), which generates motions corresponding to natural language descriptions; scene-conditioned motion generation (Araújo et al., 2023; Ghosh et al., 2023; Huang et al., 2023; Wang et al., 2022), which produces plausible human motions in relation to the surrounding scene layout; and audio-conditioned motion generation, which includes speech-to-gesture generation for conversational scenarios (Yang et al., 2023a; Alexanderson et al., 2023; Ao et al., 2023) as well as music-to-dance generation, which will be discussed in detail later.
Among these, text-conditioned motion generation has received growing attention due to the expressive power and clarity of natural language. In particular, diffusion-based approaches have recently achieved state-of-the-art performance in this area. For example, MotionDiffuse(Zhang et al., 2022) demonstrates that free-form textual input can guide the generation of diverse, coherent, and stylistically rich motion sequences. Extensions such as FLAME(Kim et al., 2023), BEing-Diffusion(Lin et al., 2023), and T2M-GEN(Zhang et al., 2023) incorporate fine-grained control over motion categories, temporal structure, and global style. UDE (Zhou and Wang, 2023) further introduces a unified representation to support cross-task generalization and improved semantic grounding. These methods demonstrate the effectiveness of using natural language as a precise, interpretable, and semantically expressive modality for guiding human motion generation.
2.2. Music-to-Dance Generation
The Music-to-Dance generation task poses unique challenges due to the need for models to understand and generate dance sequences that are both rhythmically and stylistically aligned with music. Previous methods (Chan et al., 2019; Zhuang et al., 2022; Li et al., 2023; Tseng et al., 2023; Yang et al., 2023b; Zhang et al., 2024b; Huang et al., 2024) incorporate advanced audio-processing techniques with motion generation algorithms to enhance the flexibility and expressiveness of generated dances. EDGE (Tseng et al., 2023) employs Jukebox for enriched music audio representation and introduces a contact consistency loss to reduce physical implausibilities.
Despite these advancements, existing approaches still face limitations in bridging the substantial semantic gap between music and dance, making the synchronization and aesthetic alignment of dance sequences a significant challenge. Unlike text-to-motion generation (Zhang et al., 2024a; Zhong et al., 2023), where textual cues explicitly describe body movements, music provides only implicit cues through rhythm, tempo, and tone, which do not directly correspond to physical actions. This gap complicates the creation of dance sequences that are both well-aligned with the music and stylistically diverse.
2.3. Large Language Models
Recent instruction-tuned LLMs, such as ChatGPT (OpenAI, 2023), Claude (Anthropic, 2023), LLaMA (Touvron et al., 2023), and DeepSeek (DeepSeek, 2024) have demonstrated strong capabilities in capturing contextual semantics, following user intent, and generalizing across domains. Inspired by these advancements, researchers have begun exploring LLMs for non-linguistic tasks, including motion generation (Zhou et al., 2024; Ribeiro-Gomes et al., 2024), where their ability to model temporal coherence and semantic structure enables fine-grained motion control.
To effectively integrate LLMs into non-linguistic domains, prompt engineering has emerged as a key technique (Liu et al., 2021; Liu and Chilton, 2022; Oppenlaender, 2023). This involves crafting structured prompts that effectively communicate with models, directing the generation toward desired outcomes, and allowing users to control both the content and style of the outputs. In motion generation, prompt engineering enables models to translate high-level textual instructions into semantically coherent motion sequences, enhancing controllability and expressiveness (Wang et al., 2024). Building upon these developments, our work introduces a novel approach that leverages LLM-generated textual choreography prompts to bridge the semantic gap between music and dance.
3. Our Approach
3.1. Problem Definition
The task of music-to-dance generation involves producing a dance motion sequence $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ from a given music track $\mathcal{M}$ with a duration of $\mathit{N}$ frames. Here, $\mathbf{x}_{i}\in\mathbb{R}^{D}$ represents a human pose at the i-th frame, which is denoted as a D-dimensional vector. Following the pose representation in EDGE (Tseng et al., 2023), we present the dance sequences with 24-joint SMPL format (Loper et al., 2023), which represents each joint with 6-DOF rotation representation (Zhou et al., 2019) and a single root translation. Along with the binary contact labels $b\in\{0,1\}^{2\times 2=4}$ for the heel and the toe of each foot, the dimension $D=24\times 6+3+4=151$.
3.2. Overview
Our model architecture is illustrated in Fig. 2 and consists of three main components: Pseudo Instruction Generation, Multi-modal Feature Extraction and Fusion, and Diffusion-based Motion Synthesis. The Pseudo Instruction Generation module (Fig. 2a) extracts music caption (tempo, beat, key, chord) from the given music condition $\mathcal{M}$ and utilizes an LLM as a pseudo-choreographer to transform this information into dance motion instructions $\mathcal{T}$, providing initial guidance for motion generation. The Multi-modal Fusion module (Fig. 2b) then integrates the encoded representations of music $\mathcal{M}$, beats $\mathcal{B}$, and text $\mathcal{T}$ into a unified conditional embedding. This fused representation aligns music, beat, and movement intentions, creating a cohesive context for dance generation. Finally, the Diffusion-based Motion Synthesis component (Fig. 2c) uses this unified embedding to guide the generation of realistic, rhythmically aligned dance movements that capture the style and dynamics of the music, producing smooth and expressive dance sequences.
3.3. Pseudo Instruction Generation
Figure 3. Details of our textual instruction generation process. We construct prompts using a templated format. The prompt guides an LLM to generate HumanML3D-style dance instructions with constraints on motion focus, perspective, and length. The bottom section shows an example of a prompt and its corresponding LLM-generated instruction.
This module is responsible for transforming music into high-level dance instructions, which serve as semantic guidance for the dance generation process. This module comprises two key steps: Music Caption Construction and Textual Instruction Generation.
Music Caption Construction. We extract essential musical attributes from the input music $\mathcal{M}$ to create a comprehensive music caption. Following Mustango (Melechovsky et al., 2023), we utilize multiple specialized tools to carry out the extraction:
  • Tempo: We employ BeatNet (Heydari et al., 2021) to extract tempo information, measured in beats per minute (BPM), to capture the pace of the music. This tempo provides a foundational cue for the timing of generated dance movements.
  • Chord Progression: Using Chordino (Mauch and Dixon, 2010), we identify the sequence of chords within the music. Chord progression conveys harmonic changes, which influence the emotional tone of the dance.
  • Key: The key of the music is extracted using Essentia’s (Bogdanov et al., 2013) KeyExtractor, which determines the tonal center of the piece, providing coherence and structure to the generated dance.
  • Beat Structure: BeatNet (Heydari et al., 2021) is further utilized to detect beats and downbeats, which serve as rhythmic anchors for synchronizing dance movements with the music.
Textual Instruction Generation. Given the constructed music caption, we generate a textual dance instruction $\mathcal{T}$ that aligns with the corresponding music clip $\mathcal{M}$. The instruction serves as a pseudo-guidance for training and is produced using a structured prompt directed to an LLM, which acts as a pseudo choreographer. To ensure that the generated instruction is rhythmically and stylistically appropriate, we adopt a carefully designed prompt engineering strategy inspired by the format and motion description style of the HumanML3D dataset (Guo et al., 2022), a comprehensive collection comprising diverse text-motion pair data. As shown in Fig 3, we designed a prompt engineering strategy that incorporates several key considerations to guide the model’s output effectively. First, we include the duration of the music clip in the prompt, ensuring that generated instructions are aligned with the length and pacing of the music. Additionally, to produce genre-specific movement, the prompt specifies a dancing style (e.g., hip-pop, ballet) extracted from data labels, providing the model with style supervision to create dance instructions that match both the music’s rhythm and genre. To reduce noise in the generated text, we explicitly direct the model to avoid directional words (e.g., left, right), which can introduce ambiguity or irrelevant details. This focus on general body movements without spatial directions helps maintain clarity and relevance in the instructions. We also aim to control the text style, structuring the prompt concisely and descriptively.
Instruction Results. The output of this process is a set of natural language dance instructions that provide semantic cues for the dance generation model. For instance, given a music track with a moderate tempo and a hip-hop style, the generated instruction might read: The dancer bounces lightly to the beat, alternating shoulder pops with a subtle chest pump. These instructions capture both the rhythmic alignment and stylistic characteristics of the music, bridging the semantic gap between audio and motion. Example is shown in Fig. 3.
3.4. Multi-modal Feature Extraction and Fusion
This module integrates multiple conditions from text, music, and beat information into a unified conditional embedding. We first encode each modality individually, then hierarchically fuse these embeddings to leverage their complementary strengths effectively.
3.4.1. Multi-modal Feature Extraction
Music Encoding. We leverage a pre-trained Jukebox (Dhariwal et al., 2020) model to extract high-level musical features from raw audio inputs. Specifically, we input the raw audio into the frozen Jukebox encoder, obtaining embeddings that represent the musical content. To adapt these embeddings for our task, we introduce a trainable two-layer Transformer encoder that processes the Jukebox-derived features, allowing the model to learn task-specific representations while retaining the rich musical information captured by Jukebox, resulting in embeddings $E_{M}\in\mathbb{R}^{d_{M}}$.
Text Encoding. The dance instruction generated by LLMs is processed through a CLIP (Radford et al., 2021) text encoder to capture semantic details about body movements and style. To tailor the text embeddings to our specific application, we incorporate a learnable 4-layer Transformer as a conditioning encoder. This additional Transformer processes the frozen CLIP embeddings, enabling the model to capture extra semantic dance guidance. To better leverage prior knowledge from existing text-to-motion models, we initialize our entire text encoder with the pretrained weights from MotionDiffuse (Zhang et al., 2022). This is motivated by the fact that our pseudo instructions are generated to closely follow the style and vocabulary of the HumanML3D (Guo et al., 2022) dataset, which was also used to train MotionDiffuse. Since our textual instructions are automatically generated and no paired text-motion supervision is available, we do not apply direct supervised training to the text encoder. Instead, we rely on the strong prior learned from MotionDiffuse to provide a semantically meaningful representation. The pretrained CLIP and Transformer modules together form our text encoder, and the final textual embedding is denoted as $E_{T}\in\mathbb{R}^{d_{T}}$.
Beat Encoding. In addition to basic musical features, we explicitly incorporate beat proximity information to enhance rhythmic awareness in generated motion. We apply Librosa (McFee et al., 2015) to extract both beat and downbeat positions from the audio. These positions are discrete time points that mark the perceived rhythmic structure of the music. To represent rhythmic proximity in a continuous form, we follow a Gaussian-based approach: for each frame, we compute a Gaussian-like pulse centered around the nearest beat and downbeat positions. This encoding reflects the temporal distance to nearby rhythmic accents, enabling the model to perceive and respond to the rhythmic context over time. We then project these proximity values into a learnable embedding space using a linear projection layer, resulting in beat embeddings $E_{B}\in\mathbb{R}^{d_{B}}$.
Together, these extracted features capture the unique attributes of each modality, providing a foundation for effective cross-modal integration. Each of these extracted feature embeddings provides distinct complementary information: textual embedding $E_{T}$ offers semantic pseudo guidance, music embedding $E_{M}$ captures musical style, and beat embedding $E_{B}$ encodes rhythmic timing.
3.4.2. Multi-modal Hierarchical Fusion
We fuse these modality embeddings through a hierarchical, two-level strategy designed to incrementally integrate rhythmic, musical, and textual guidance.
Musical-Rhythmic Fusion. Initially, we fuse music embeddings $E_{M}$ with rhythmic beat embedding $E_{B}$ using an additive fusion approach:
(1) $E_{M}^{\prime}=E_{M}+E_{B},$
This straightforward yet effective fusion explicitly integrates rhythmic timing information directly into the music representations. By enhancing music embeddings with beat cues, it enables the model to accurately capture rhythmically significant moments within the music.
Music-Text Semantic Fusion. Subsequently, we incorporate textual information from dance instruction embeddings $E_{T}$ to provide semantic guidance for dance generation. Specifically, we concatenate the textual embeddings $E_{T}$ with the previously fused musical-rhythmic embeddings $E_{M}^{\prime}$:
(2) $E_{F}=[E_{M}^{\prime};E_{T}],$
generating the final unified conditional embedding $E_{F}$. This second fusion stage integrates high-level choreographic instructions, derived from natural language guidance, with detailed rhythmic-musical context. Thus, the final embedding $E_{F}$ effectively combines semantic choreographic concepts from textual instructions and detailed rhythmic structure from music and beat embeddings.
The final conditional embedding $E_{F}$ encapsulates a balanced representation of the text, music, and beat, ensuring that generated dance movements are not only synchronized with the rhythm but also stylistically aligned and semantically meaningful. This fused embedding serves as the input for the diffusion-based motion synthesis module, enabling the generation of realistic, expressive dance sequences that reflect the musical nuances and choreographic intentions embedded in the text.
3.5. Diffusion-based Motion Synthesis
This module leverages a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) to progressively guide noisy samples toward realistic dance motions, conditioned on the multi-modal embedding $E_{F}$ from the previous Multi-modal Fusion stage.
We model dance generation as a Markovian diffusion process, in which a sequence of noisy latent representations is gradually denoised to reveal the desired motion. The forward diffusion process begins with a clean motion sequence $\mathcal{X}$ and iteratively adds Gaussian noise over $T$ timesteps, eventually transforming it into a nearly random noise sample $z_{T}$. Formally, at each timestep $t$, the forward process is defined as:
(3) $\mathit{q(z_{t}|z_{t-1})}=\mathcal{N}(\sqrt{\mathit{1-\beta_{t}}}\mathit{z_{t-1}},\beta_{t}\mathit{I}),$
where $\beta_{t}\in(0,1)$ is a variance schedule parameter that controls the amount of noise added at each step. By the final timestep $T$, the noisy sequence $z_{T}$ closely approximates a standard Gaussian distribution $\mathcal{N}(0,\mathit{I})$.
To recover the original motion from this noisy representation, we employ a reverse diffusion process. This reverse process begins with the noisy sample $z_{T}$ and iteratively denoises it, guided by the conditional embedding $E_{F}$, to approximate the original sequence $\mathcal{X}$. At each step $t$, we predict the previous timestep $z_{t-1}$ from $z_{t}$ using:
(4) $\mathit{p_{\theta}(z_{t-1}|z_{t},E_{F})}=\mathcal{N}(\mathit{z_{t-1}};\mu_{\theta}(z_{t},t,E_{F}),\sigma_{t}^{2}\mathit{I}),$
where $\mu_{\theta}(z_{t},t,E_{F})$ represents the predicted mean for $z_{t-1}$ based on the model parameters $\theta$, and $\sigma_{t}$ is a learned variance that adjusts the noise schedule during the denoising process.
We adopt the simplified training objective from DDPM (Ho et al., 2020), minimizing the mean squared error (MSE) between the model’s predictions and the true motion data across all timesteps. This objective encourages the model to reconstruct realistic dance sequences from noisy inputs while aligning with the semantic and rhythmic cues encoded in $E_{F}$. Formally, the training objective is defined as:
(5) $\mathcal{L}_{\mathrm{simple}}=\mathbb{E}_{\mathbf{x},t}[\parallel\mathbf{x}-\mathbf{\hat{x}}(z_{t},t,E_{F})\parallel^{2}_{2}],$
where $\mathbf{\hat{x}}(z_{t},t,E_{F})$ denotes the reconstructed motion at timestep $t$ conditioned on the multi-modal embedding $E_{F}$.
3.6. Loss Functions
Method PFC $\downarrow$ $\mathrm{Div}_{k}\rightarrow$ $\mathrm{Div}_{g}\rightarrow$ Beat Align. Score $\uparrow$
Ground Truth 1.332 10.61 7.48 0.24
FACT (Li et al., 2021) 2.254 10.85 6.14 0.22
Bailando (Siyao et al., 2022) 1.754 7.92 7.72 0.23
EDGE (Tseng et al., 2023) 1.536 9.48 5.72 0.23
BADM (Zhang et al., 2024b) 1.424 8.29 6.76 0.24
Beat-it (Huang et al., 2024) 0.966 9.66 7.25 0.66
Ours DanceChat 0.828 10.64 7.36 0.27
Table 1. Quantitative comparison with state-of-the-art methods on the AIST++ dataset. FACT and Bailando are Transformer- and VQ-VAE-based models that require a starting pose. EDGE, BADM, and Beat-it are diffusion-based approaches. $\uparrow$ means higher is better, $\downarrow$ means lower is better, $\rightarrow$ means closer to ground truth is better. Bold indicated the best result; underline indicates the second-best. Our method outperforms the other methods on $\mathrm{PFC}$, $\mathrm{DIV}_{k}$ and $\mathrm{DIV}_{g}$, and performs the second-best on $\mathrm{BAS}$.
Kinematic Loss. To enhance the physical plausibility of generated motions, we incorporate several auxiliary losses commonly used in kinematic motion generation. In addition to the basic reconstruction loss $\mathcal{L}_{\mathrm{simple}}$, we adopt auxiliary terms introduced by Tevet et al.(Tevet et al., 2022b) and further refined in EDGE(Tseng et al., 2023), targeting three key aspects of motion realism: joint positions (Eq. 6), joint velocities (Eq. 7), and foot-ground contact consistency (Eq. 8). Below, $\mathbf{FK(\cdot)}$ denotes the forward kinematic function that converts joint angles into joint positions and the $\mathit{i}$-th superscript denotes the frame index. We note Eq. 8 only applies to the foot joints, and $\mathbf{\hat{b}}_{i}$ is binary foot contact signal.
(6) $\mathcal{L}_{\mathrm{joint}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}\parallel\mathbf{FK}(\mathbf{x}_{i})-\mathbf{FK}(\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$
(7) $\mathcal{L}_{\mathrm{vel}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{x}_{i+1}-\mathbf{x}_{i})-(\mathbf{\hat{x}}_{i+1}-\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$
(8) $\mathcal{L}_{\mathrm{contact}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{FK}(\mathbf{x}_{i+1})-\mathbf{FK}(\mathbf{\hat{x}}_{i}))\cdot\mathbf{\hat{b}}_{i}\parallel^{2}_{2}$
Multi-modal Alignment Loss. To bridge the semantic gap between music and motion, we design a multi-modal alignment loss function that leverages text as an intermediate modality. Unlike traditional approaches that directly align music and motion, our method calculates similarity between motion-text and text-music pairs at each timestep $\mathit{t}$ in the diffusion process. By using text as a common semantic space, we allow the model to capture descriptive and interpretable features that relate both to the audio’s rhythmic and tonal characteristics and to the motion’s dynamic and stylistic attributes. Specifically, for each timestep, we compute cosine similarity between motion and text embeddings and text and music embeddings, promoting alignment across all three modalities through the text-based intermediary. The total alignment loss is defined as Eq. 9, encouraging consistent semantic alignment between music and motion via the shared text representation.
(9) $\mathcal{L}_{\mathrm{align}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}[(1-\mathrm{sim}(E_{t_{i}},E_{m_{i}}))+(1-\mathrm{sim}(E_{t_{i}},E_{x_{i}}))]$
This alignment loss is then combined with the other auxiliary losses, weighted by factors $\lambda$ to balance the magnitudes of the losses at the start of training.
(10) $\begin{split}\mathcal{L}&=\mathcal{L}_{\mathrm{simple}}+\lambda_{\mathrm{joint}}\mathcal{L}_{\mathrm{joint}}+\lambda_{\mathrm{vel}}\mathcal{L}_{\mathrm{vel}}+\\ &\lambda_{\mathrm{contact}}\mathcal{L}_{\mathrm{contact}}+\lambda_{\mathrm{align}}\mathcal{L}_{\mathrm{align}}\end{split}$
4. Experiments
4.1. Experiment Setup
Figure 4. Example results generated by our model. Each sequence illustrates a 3D dance motion generated from a given music clip, conditioned on an LLM-generated textual instruction. The sentences below each motion sequence are produced by the LLM and serve as choreography prompts.
Dataset. We use the AIST++(Li et al., 2021) dataset for both training and testing. This dataset comprises 1,408 dance sequences paired with music, covering 10 distinct street dance genres. Each genre is synchronized with music from 6 different tracks, each with unique beats per minute (BPM) and musical characteristics, allowing for a variety of rhythmic and stylistic alignment in the generated dance motions. All dance motions are represented as 3D poses captured at 60 FPS, providing detailed joint-level information for each frame. For our experiments, we segment these music-motion pair sequences into 5-second clips sampled at 30 FPS, and extract music caption and textual guidance based on each segmented music clip. We follow the standard train/test splits provided in the original dataset.
Implementation details. In the pseudo instruction generation module, we utilize the OpenAI ChatGPT-4o API (OpenAI API, [n. d.]) to generate textual dance motion instruction based on pre-extracted music captions. For prompt engineering, we use the annotated genre information during training. At test time, we replace the genre placeholder in the prompt with the literal word any. In the multi-modal feature extraction and fusion module, we adopt an encoding approach in which pre-trained models are frozen to preserve general representations, while lightweight trainable layers are added for task-specific adaptation. We first extract features using a frozen Jukebox model and subsequently apply two additional Transformer encoder layers to refine the temporal and semantic representations. For text encoding, we use the frozen text encoder from CLIP ViT-B/32 (Radford et al., 2021) followed by another two-layer Transformer encoder to adapt the output for conditional guidance. The latent dimensionality of both the music and text encoders is set to 512. Our model has 54M total parameters, and was trained on 1 NVIDIA A100 GPU for 48 hours with a batch size of 128.
4.2. Comparison to Existing Methods
We have selected FACT (Li et al., 2021), Bailando (Siyao et al., 2022), EDGE (Tseng et al., 2023), BADM (Zhang et al., 2024b) and Beat-it (Huang et al., 2024) for comparisons with our DanceChat method. FACT utilizes a full-attention mechanism to capture dependencies across music and motion, leveraging autoregressive generation to maintain consistency in dance sequences. Bailando builds upon this by incorporating a VQ-VAE (Van Den Oord et al., 2017) and GPT (Radford, 2018) framework, enhanced with an actor-critic reinforcement learning scheme to improve dance quality. EDGE introduces a transformer-based diffusion model, serving as the backbone of our method. BADM, as a follow-up to EDGE, proposes a bidirectional autoregressive diffusion model that enforces harmony in both forward and backward directions, capturing local motion details for smoother dance sequences. Similarly, Beat-it extends EDGE by presenting a beat-synchronized, key pose-guided 3D dance generation framework, integrating explicit beat awareness and key pose guidance to ensure alignment between dance motions and musical beats.
To evaluate the effectiveness of our approach, we conduct a series of experiments to assess the performance in terms of physical plausibility, motion diversity, and beat alignment score (BAS), which together evaluate the physical realism, stylistic variability, and rhythmic quality of the generated dance motion. The quantitative metrics are shown in Table 1. In addition to this, we also conduct a human evaluation to gather subjective feedback on the overall quality and coherence of the generated motions, as shown in Table 2.
Motion Quality. We incorporate the Physical Foot Contact score (PFC) (Tseng et al., 2023) to evaluate the foot-ground interactions, a critical component of physical realism in dance generation. The PFC score assesses the accuracy of foot contact behavior, ensuring that generated sequences avoid unrealistic foot sliding and maintain natural center of mass (COM) dynamics. A lower PFC score indicates that the generated movements more closely approximate realistic foot-ground interactions, with stable foot placements and controlled COM dynamics, leading to improved balance and physical plausibility. As shown in Table 1, DanceChat achieves the best performance in $\mathrm{PFC}$, with an improvement of 0.138 compared with SOTA method Beat-it (Huang et al., 2024). This result highlights DanceChat’s ability to generate dance sequences with high physically realistic foot-ground interactions.
Motion Diversity. To assess the diversity of the generated dance motions, we follow the approach outlined by Bailando (Siyao et al., 2022). We compute the mean Euclidean distance within the motion feature space to quantify the spread of generated motions. Specifically, we calculate two diversity metrics: $\mathrm{Div}_{k}$, which reflects diversity in the kinematic feature space, and $\mathrm{Div}_{g}$, which measures diversity in the geometric feature space. These metrics are computed over 5-second dance clips to maintain consistency with prior benchmarks (Tseng et al., 2023; Zhang et al., 2024b; Huang et al., 2024). Following observations from EDGE (Tseng et al., 2023), excessively jittery or unstable motions can lead to artificially inflated diversity scores, occasionally exceeding those of the ground truth. Therefore, rather than maximizing diversity, models should aim to match the distributional spread of the ground truth, indicating an appropriate balance between variation and realism. As shown in Table 1, our approach achieves a $\mathrm{Div}_{g}$ score of 7.36 and a $\mathrm{Div}_{k}$ score of 10.64, which are closer to the ground truth than all the other methods. We attribute to our model’s use of global diffusion techniques and fundamental dance primitives, enabling it to learn diverse choreography patterns.
Beat Alignment Score (BAS). To evaluate the rhythmic alignment of generated dance movements with the accompanying music, we use the BAS metric, which measures the consistency between the music beats and the kinematic beats of the generated motion. Following Bailando (Siyao et al., 2022), this metric rewards alignment when local minima in joint speed (indicative of pauses or transitions in motion) correspond with the musical beats. As shown in Table 1, our model achieves the second-best BAS score, indicating a strong ability to follow the beat and maintain rhythmic coherence. The highest BAS is achieved by Beat-it (Huang et al., 2024), which explicitly optimizes for beat alignment using a strong beat-supervised loss combined with key-pose constraints, allowing it to tightly synchronize motions with musical beats. In contrast, our model achieves competitive rhythmic alignment without requiring strong beat-level supervision. Instead, we rely on a flexible multi-modal fusion framework that incorporates both musical and beat features in a weakly supervised manner. This enables our model to produce rhythmically aware motion while maintaining better generalization across genres and motion styles.
Human Evaluation. To gain a deeper understanding of the visual and rhythmic quality of our method, we conduct a user study comparing the dance sequences generated by DanceChat with those produced by baseline methods on the AIST++ dataset and the in-the-wild musics. This study involves 30 participants, each engaging in individual evaluations. Participants are shown a series of 20 pairs of comparison videos, each approximately 10 seconds in length. Each pair includes one video generated by DanceChat and one by a competing method, synchronized to the same music track. The participants’ task is to identify which of the two videos feel better. The results of the user study on AIST++ are summarized in Table 2. Notably, DanceChat outperformed all the baseline results. We also conduct a separate evaluation in an “in-the-wild” setting, using unseen music tracks to test DanceChat’s generalization capabilities. As shown in Table 3, DanceChat achieves a preference rate of 66.36% over EDGE in these real-world scenarios, further confirming its robustness and adaptability across different musical contexts. In addition to the user study, an observable reduction in foot sliding is noted in the DanceChat videos, further confirming the PFC improvement.
4.3. Ablation Study
We perform an ablation study on the AIST++ dataset to evaluate the impact of different modalities (music, beat, and text) and the effectiveness of the multi-modal alignment loss in generating high-quality dance motions. The results, shown in Table 4, are reported in terms of $\mathrm{PFC}$ (for motion realism), $\mathrm{DIV}_{k}$ (for motion diversity), and BAS (for beat alignment score), with lower $\mathrm{PFC}$ and higher $\mathrm{DIV}_{k}$ and BAS indicating better performance.
Impact of Modality Inputs. To understand the individual contributions of each modality, we assess various configurations by selectively including or excluding the music (M), beat (B), and text (T) modalities. The baseline model, using only music (M), performs the lowest across all metrics, with limited diversity and weak rhythmic alignment. Adding beat information (M+B) enhances rhythmic alignment, indicating that beat cues contribute both to physical realism in generated movement patterns and to synchronization with the music. Adding textual guidance (M+T) leads to the best performance in PFC and diversity, suggesting that text improves motion realism and expressiveness by providing explicit semantic choreography cues. However, lacking beat information, this configuration underperforms in BAS. Interestingly, combining all three modalities (M+B+T) produces the highest BAS and consistently strong results in other metrics.
Impact of Multi-Modal Alignment Loss. We further incorporate the multi-modal alignment loss (MM Align. Loss), which leverages text as an intermediary bridge between music and motion to narrow the semantic gap between these modalities. The alignment loss leads to both improved $\mathrm{PFC}$) and $\mathrm{DIV}_{k}$, demonstrating its effectiveness in facilitating cross-modal integration. This suggests that using text as a bridge in the alignment loss helps create semantically rich and rhythmically aligned dance motions that benefit from the strengths of each modality without compromising diversity or quality.
[…]
FACT (Li et al., 2021) 86.25%
Bailando (Siyao et al., 2022) 61.88%
EDGE (Tseng et al., 2023) 60.00%
Ground Truth 48.12%
Table 2. User study results on AIST++ test set. We present the percentage of user evaluations that prefer DanceChat to FACT (Li et al., 2021), Bailando (Siyao et al., 2022), and EDGE (Tseng et al., 2023).
[…]
EDGE (Tseng et al., 2023) 66.36%
Table 3. User study results of in-the-wild music. We compare DanceChat against EDGE(Tseng et al., 2023).
Ablations Metrics
M B T PFC $\downarrow$ Divk $\rightarrow$ Beat Align. Score $\uparrow$
1.536 9.48 0.26
1.066 8.76 0.27
1.166 9.69 0.26
0.843 9.85 0.28
MM Align. Loss 0.828 10.64 0.27
Ground Truth 1.332 10.61 0.24
Table 4. Ablation study of modalities, where M indicates music, B indicates beat and T indicates text. The best results are indicated in bold, and the second best are underlined.
5. Conclusion
In this paper, we present DanceChat, a multi-modal, language-guided framework for music-to-dance generation, addressing the significant semantic gap between music and dance. By leveraging a Large Language Model (LLM) as a pseudo-choreographer, our approach introduces natural language as an intermediary between music and motion, transforming abstract musical elements into concrete dance instructions. Our proposed model architecture comprises three core components: Pseudo Instruction Generation, which translates musical meta-information into text-based dance instructions; Multi-modal Fusion, which integrates music, beat, and textual guidance into a cohesive representation; and Diffusion-based Motion Synthesis, which employs a denoising process to generate smooth, expressive dance motions. Through comprehensive experiments on the AIST++ dataset, DanceChat demonstrates competitive performance across multiple metrics. The results confirm that each modality contributes uniquely to dance quality, while our multi-modal alignment loss plays a crucial role in harmonizing these modalities to produce musically aligned sequences.
Limitations and Future Work. While DanceChat achieves strong results, certain limitations remain. First, the reliance on pre-trained LLMs and external encoders introduces a dependency on existing models that may not be fully optimized for dance generation. Future work could explore custom training for language models specifically tailored for dance, potentially enhancing the semantic richness of generated instructions.

### v2

@@
 DanceChat: Large Language Model-Guided Music-to-Dance Generation
-Abstract.
-Music-to-dance generation aims to synthesize human dance motion conditioned on musical input.
-Despite recent progress, significant challenges remain due to the semantic gap between music and dance motion, as music offers only abstract cues, such as melody, groove, and emotion, without explicitly specifying the physical movements.
-Moreover, a single piece of music can produce multiple plausible dance interpretations.
-This one-to-many mapping demands additional guidance, as music alone provides limited information for generating diverse dance movements.
+Abstract
+Music-to-dance generation aims to synthesize human dance motion conditioned on music input.
+Despite recent progress, significant challenges remain due to the semantic gap between music and dance motion, as music offers only abstract cues, but lacks explicit physical movement descriptions.
 The challenge is further amplified by the scarcity of paired music and dance data, which restricts the model’s ability to learn diverse dance patterns.
-In this paper, we introduce DanceChat, a Large Language Model (LLM)-guided music-to-dance generation approach.
-We use an LLM as a choreographer that provides textual motion instructions, offering explicit, high-level guidance for dance generation.
-This approach goes beyond implicit learning from music alone, enabling the model to generate dance that is both more diverse and better aligned with musical styles.
-Our approach consists of three components: (1) an LLM-based pseudo instruction generation module that produces textual dance guidance based on music style and structure, (2) a multi-modal feature extraction and fusion module that integrates music, rhythm, and textual guidance into a shared representation, and (3) a diffusion-based motion synthesis module together with a multi-modal alignment loss, which ensures that the generated dance is aligned with both musical and textual cues.
-Extensive experiments on AIST++ and human evaluations show that DanceChat outperforms state-of-the-art methods both qualitatively and quantitatively.
-The additional qualitative samples can be found at https://dancechat.github.io/anon/.
-3D Motion Synthesis, Cross-Modality Learning, Tempo Synchronization
-ccs: Computing methodologies Motion processingccs: Applied computing Sound and music computing
-Figure 1.
-Our approach inspired by real-world dance learning.
-Just as dance learners rely on choreographers to interpret music, DanceChat uses textual dance instructions to bridge the semantic gap between music and motion.
-Large Language Models (LLMs) act as pseudo-choreographers, transforming musical cues into detailed, actionable guidance for dance generation.
-1. Introduction
-Dance has long served as a powerful medium of expression in cultures, playing a central role in rituals, social connections, and artistic activities (Carroll and Moore, 2008).
+These limitations highlight the need for additional semantic guidance beyond the musical signal.
+In this paper, we propose DanceChat, a novel framework that leverages a Large Language Model (LLM) as a choreographer to generate high-level textual instructions from structured music descriptions.
+These instructions serve as semantic guidance to bridge the gap between music and motion.
+DanceChat integrates music, beat, and text features into a unified representation, and employs a diffusion-based motion generator trained with a proposed multi-modal alignment loss.
+Extensive experiments on AIST++ dataset show that DanceChat outperforms state-of-the-art methods both qualitatively and quantitatively.
+1 Introduction
+Dance has long served as a powerful medium of expression in cultures, playing a central role in rituals, social connections, and artistic activities (Carroll and Moore 2008).
 In the digital age, dance flourishes on social media, allowing people to share creativity and connect globally.
 Recent advances in deep learning have enabled computational exploration of this art form, particularly through music-to-dance generation, where models automatically create dance movements synchronized to music.
-This technology has impactful applications in entertainment, enabling realistic, rhythmically aligned virtual idols and 3D characters in virtual reality.
-Beyond entertainment, automated dance generation is valuable in fields like robotics, digital human modeling, and physical rehabilitation, where it helps create lifelike movement simulations and supports human-computer interaction.
-Most existing music-to-dance generation approaches fall into two main categories: regression models (Huang et al., 2021; Li et al., 2021), which are typically designed to directly learn a mapping from input music to dance motions, and generative models (Siyao et al., 2022; Tseng et al., 2023; Zhang et al., 2024b; Huang et al., 2024), which aim to learn a distribution over plausible dances conditioned solely on the input music.
-However, both approaches lack explicit movement instructions, and rely only on abstract nature of musical elements such as rhythm, tempo, and mood.
-This creates a broad semantic gap between music and motion, as illustrated in Fig.
-1, and complicates the task of generating physically plausible movements that align with the music’s genre and maintain aesthetic quality.
+This technology has impactful applications in entertainment, enabling realistic virtual idols and 3D characters in virtual reality.
+Beyond entertainment, automated dance generation is valuable in fields like robotics, digital human, and physical rehabilitation, where it helps create lifelike movement simulations and supports human-computer interaction.
+Figure 1: Our approach draws inspiration from real-world dance learning, where learners rely on choreographers to interpret music.
+DanceChat leverages LLMs as pseudo-choreographers to translate music into textual instructions, bridging the semantic gap between music and motion.
+While the mapping from music to motion is inherently one-to-many and ill-posed, text offers a more structured and interpretable intermediary.
+Most existing music-to-dance generation approaches fall into two main categories: regression models (Huang et al.
+2021; Li et al.
+2021), which are typically designed to directly learn a mapping from input music to dance motions, and generative models (Siyao et al.
+2022; Tseng, Castellon, and Liu 2023; Zhang et al.
+2024a; Huang et al.
+2024), which aim to learn a distribution over plausible dances conditioned solely on the input music.
+However, both approaches lack explicit movement instructions, and rely only on the abstract nature of musical elements.
+This creates a broad semantic gap between music and motion and complicates the task of generating physically plausible movements that align with the music’s genre and maintain aesthetic quality (Lee et al.
+2019).
 Moreover, the scarcity of labeled dance data, due in part to the high cost of motion capture, makes it difficult for models to learn a well-generalized and diverse motion distribution.
@@
 These challenges highlight the need for advanced methods capable of bridging complex semantic gaps and injecting semantic-rich textual instructions without relying on additional annotations.
-Recent advancements in large language models (LLMs) have demonstrated remarkable success in capturing nuanced semantic relationships across various domains (Wu et al., 2023; Han et al., 2023; Doh et al., 2024; Gardner et al., 2024; Ding et al., 2024).
-They are trained on diverse content, including detailed discussions on music, enabling them to interpret musical elements such as tempo, key, and chord progressions.
-This allows them to understand and generate contextually appropriate descriptions that capture the structure, rhythm, and mood of a piece.
-This inspires new approaches for tasks involving multi-modal relationships and encourages us to investigate the possibilities for leveraging LLMs to enhance the controllability of music-to-dance generation.
+Recent advancements in large language models (LLMs) have demonstrated remarkable success in capturing semantic relationships across various domains (Wu et al.
+2023; Han et al.
+2023; Doh et al.
+2024; Gardner et al.
+2024; Ding et al.
+2024).
+They are trained on diverse content, including detailed discussions on music and dance, enabling them to interpret musical elements such as tempo, key, and chord progressions, as well as recognize movement styles, choreographic vocabulary, and dance semantics.
+This allows them to generate contextually appropriate descriptions that capture not only the structure and rhythm of a musical piece, but also align with plausible dance patterns.
+This inspires us to explore the potential of leveraging LLMs to enhance the diversity of music-to-dance generation.
 Building on these insights, our method adapts the capabilities of LLMs to address the semantic gap in music-driven dance generation.
-We propose a novel approach, named DanceChat, which conceptualizes the generation model as a dancing learner and leverages LLMs as a pseudo-choreographer.
-Specifically, we extract structured musical descriptors, including key, tempo, and chord progressions alongside the dance genre context, and feed this information into LLM.
-The model then generates corresponding textual dance instructions, which serve as a pseudo-guide to condition the motion generation model during training.
+We propose a novel approach, named DanceChat, which conceptualizes the generation model as a dancing learner and leverages LLMs as a pseudo-choreographer (Fig.
+1).
+Specifically, we extract structured musical descriptors, including key, tempo, and chord progressions alongside the dance genre context, and feed this information into LLM, which then generates corresponding textual dance instructions, serving as a pseudo-guide to condition the motion generation model during training.
 This helps reduce the semantic gap by offering interpretative cues that map abstract musical elements to concrete motion patterns.
-To further enhance alignment between modalities, we encode the input music into finer-grained descriptors such as beat and downbeat positions, which are known to correlate strongly with movement timing in dance.
-These rhythmic cues, together with the LLM-generated textual instructions, are integrated through a multi-modal fusion module, enabling the model to form a richer and more coherent understanding of the music.
-By combining musical structure, rhythmic signals, and semantic instructions, DanceChat produces dance motions that are rhythmically aligned, physically plausible, and stylistically expressive.
+To further enhance cross-modal alignment, we explicitly extract beat and downbeat information from the input music as an additional rhythmic modality, which is known to correlate strongly with movement timing in dance.
+The original music, rhythmic cues, and LLM-generated textual instructions are then integrated through a multi-modal fusion module, enabling the model to form a richer and more coherent understanding of the music.
+By combining music, beat signals, and semantic instructions, DanceChat produces dance motions that are rhythmically aligned, physically plausible, and stylistically expressive.
 In summary, our contributions are three-fold:
@@
 3) A multi-modal alignment loss that leverages text as an intermediary to enhance the semantic connection between music and dance, reducing the gap between these modalities and improving the coherence and expressiveness of generated dance movements.
-2. Related Work
-Figure 2.
-Overview of our approach.
-DanceChat consists of three main components: (a) Pseudo Instruction Generation module extracts music caption (tempo, key, chord, etc.) from the given music condition, (b) Multi-modal Fusion module integrates the encoded representations of music, beats, and text into a unified conditional embedding, and (c) Diffusion-based Motion Synthesis module uses this unified embedding to guide the generation of realistic dance movement.
-2.1. 3D Human Motion Generation
-Common approaches to 3D human motion generation (Zhu et al., 2023) can be broadly categorized into three sub-tasks: text-conditioned motion generation (Ahn et al., 2018; Tevet et al., 2022a; Chen et al., 2023; Zhang et al., 2022; Kim et al., 2023; Lin et al., 2023; Zhang et al., 2023; Zhou and Wang, 2023), which generates motions corresponding to natural language descriptions; scene-conditioned motion generation (Araújo et al., 2023; Ghosh et al., 2023; Huang et al., 2023; Wang et al., 2022), which produces plausible human motions in relation to the surrounding scene layout; and audio-conditioned motion generation, which includes speech-to-gesture generation for conversational scenarios (Yang et al., 2023a; Alexanderson et al., 2023; Ao et al., 2023) as well as music-to-dance generation, which will be discussed in detail later.
-Among these, text-conditioned motion generation has received growing attention due to the expressive power and clarity of natural language.
-In particular, diffusion-based approaches have recently achieved state-of-the-art performance in this area.
-For example, MotionDiffuse(Zhang et al., 2022) demonstrates that free-form textual input can guide the generation of diverse, coherent, and stylistically rich motion sequences.
-Extensions such as FLAME(Kim et al., 2023), BEing-Diffusion(Lin et al., 2023), and T2M-GEN(Zhang et al., 2023) incorporate fine-grained control over motion categories, temporal structure, and global style.
-UDE (Zhou and Wang, 2023) further introduces a unified representation to support cross-task generalization and improved semantic grounding.
-These methods demonstrate the effectiveness of using natural language as a precise, interpretable, and semantically expressive modality for guiding human motion generation.
-2.2. Music-to-Dance Generation
+2 Related Works
+Figure 2: Overview of our approach.
+DanceChat consists of three main components: (a) Dance Instruction Generation module extracts music caption (tempo, key, chord, etc.) from the given music condition, (b) Multi-modal Fusion module integrates the encoded representations of music, beats, and text into a unified conditional embedding, and (c) Diffusion-based Motion Synthesis module uses this unified embedding to guide the generation of realistic dance movement.
+2.1 3D Human Motion Generation
+3D human motion generation has become a central problem in computer vision and graphics, with applications in animation, robotics, and virtual agents.
+Recent research has explored a wide range of conditioning modalities, including natural language, audio, scene layout, and control signals, aiming to improve realism, diversity, and controllability.
+Among these, text-conditioned motion generation has emerged as a promising direction, leveraging the semantic clarity and expressiveness of natural language to provide interpretable and flexible guidance.
+Diffusion-based approaches (Kim, Kim, and Choi 2023; Zhang et al.
+2024b; Liang et al.
+2024; Zhou et al.
+2024) have demonstrated that structured textual inputs can guide the synthesis of diverse and coherent motion.
+This highlights the potential of language as a strong semantic prior for motion planning.
+Building on this idea, our work proposes to use language as a semantic bridge between music and motion.
+By translating musical attributes into textual instructions, we enable explicit, high-level guidance for music-to-dance generation, addressing the limitations of relying solely on music features.
+2.2 Music-to-Dance Generation
 The Music-to-Dance generation task poses unique challenges due to the need for models to understand and generate dance sequences that are both rhythmically and stylistically aligned with music.
-Previous methods (Chan et al., 2019; Zhuang et al., 2022; Li et al., 2023; Tseng et al., 2023; Yang et al., 2023b; Zhang et al., 2024b; Huang et al., 2024) incorporate advanced audio-processing techniques with motion generation algorithms to enhance the flexibility and expressiveness of generated dances.
-EDGE (Tseng et al., 2023) employs Jukebox for enriched music audio representation and introduces a contact consistency loss to reduce physical implausibilities.
-Despite these advancements, existing approaches still face limitations in bridging the substantial semantic gap between music and dance, making the synchronization and aesthetic alignment of dance sequences a significant challenge.
-Unlike text-to-motion generation (Zhang et al., 2024a; Zhong et al., 2023), where textual cues explicitly describe body movements, music provides only implicit cues through rhythm, tempo, and tone, which do not directly correspond to physical actions.
+Previous methods (Chan et al.
+2019; Zhuang et al.
+2022; Li et al.
+2023; Tseng, Castellon, and Liu 2023; Yang, Yang, and Wang 2023; Tseng, Castellon, and Liu 2023; Huang et al.
+2024) incorporate advanced audio-processing techniques with motion generation algorithms to enhance the flexibility and expressiveness of generated dances.
+Despite these advancements, existing approaches still face limitations in bridging the substantial semantic gap between music and dance, making the synchronization of dance sequences a significant challenge.
+Unlike text-to-motion generation where textual cues explicitly describe body movements, music provides only implicit cues through rhythm, tempo, and chord, which do not directly correspond to physical actions.
 This gap complicates the creation of dance sequences that are both well-aligned with the music and stylistically diverse.
-2.3. Large Language Models
-Recent instruction-tuned LLMs, such as ChatGPT (OpenAI, 2023), Claude (Anthropic, 2023), LLaMA (Touvron et al., 2023), and DeepSeek (DeepSeek, 2024) have demonstrated strong capabilities in capturing contextual semantics, following user intent, and generalizing across domains.
-Inspired by these advancements, researchers have begun exploring LLMs for non-linguistic tasks, including motion generation (Zhou et al., 2024; Ribeiro-Gomes et al., 2024), where their ability to model temporal coherence and semantic structure enables fine-grained motion control.
-To effectively integrate LLMs into non-linguistic domains, prompt engineering has emerged as a key technique (Liu et al., 2021; Liu and Chilton, 2022; Oppenlaender, 2023).
+2.3 Large Language Models
+Recent instruction-tuned LLMs (OpenAI 2023; Anthropic 2023; Touvron et al.
+2023; DeepSeek 2024) have demonstrated strong capabilities in capturing contextual semantics, following user intent, and generalizing across domains.
+Inspired by these advancements, researchers have begun exploring LLMs for non-linguistic tasks, including motion generation (Zhou, Wan, and Wang 2024; Ribeiro-Gomes et al.
+2024), where their ability to model temporal coherence and semantic structure enables fine-grained motion control.
+To effectively integrate LLMs into non-linguistic domains, prompt engineering has emerged as a key technique (Liu et al.
+2021; Liu and Chilton 2022; Oppenlaender 2023).
 This involves crafting structured prompts that effectively communicate with models, directing the generation toward desired outcomes, and allowing users to control both the content and style of the outputs.
-In motion generation, prompt engineering enables models to translate high-level textual instructions into semantically coherent motion sequences, enhancing controllability and expressiveness (Wang et al., 2024).
-Building upon these developments, our work introduces a novel approach that leverages LLM-generated textual choreography prompts to bridge the semantic gap between music and dance.
-3. Our Approach
-3.1. Problem Definition
-The task of music-to-dance generation involves producing a dance motion sequence $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ from a given music track $\mathcal{M}$ with a duration of $\mathit{N}$ frames.
-Here, $\mathbf{x}_{i}\in\mathbb{R}^{D}$ represents a human pose at the i-th frame, which is denoted as a D-dimensional vector.
-Following the pose representation in EDGE (Tseng et al., 2023), we present the dance sequences with 24-joint SMPL format (Loper et al., 2023), which represents each joint with 6-DOF rotation representation (Zhou et al., 2019) and a single root translation.
+In motion generation, prompt engineering enables models to translate high-level textual instructions into semantically coherent motion sequences, enhancing controllability and expressiveness (Wang et al.
+2024).
+Building upon these developments, our work introduces a novel approach that leverages LLM-generated textual choreography instructions to bridge the semantic gap between music and dance.
+3 Our Approach
+3.1 Problem Definition
+The task of music-to-dance generation involves producing a dance motion sequence $\mathcal{X}=\{\mathit{x}^{1},\mathit{x}^{2},\dots,\mathit{x}^{N}\}$ from a given music track $\mathcal{M}$ with a duration of $\mathit{N}$ frames.
+Here, $\mathit{x}^{i}\in\mathbb{R}^{D}$ represents a human pose at the i-th frame, which is denoted as a D-dimensional vector.
+We present the dance sequences with 24-joint SMPL format (Loper et al.
+2023), which represents each joint with 6-DOF rotation representation (Zhou et al.
+2019) and a single root translation.
 Along with the binary contact labels $b\in\{0,1\}^{2\times 2=4}$ for the heel and the toe of each foot, the dimension $D=24\times 6+3+4=151$.
-3.2. Overview
+3.2 Overview
 Our model architecture is illustrated in Fig.
-2 and consists of three main components: Pseudo Instruction Generation, Multi-modal Feature Extraction and Fusion, and Diffusion-based Motion Synthesis.
-The Pseudo Instruction Generation module (Fig.
-2a) extracts music caption (tempo, beat, key, chord) from the given music condition $\mathcal{M}$ and utilizes an LLM as a pseudo-choreographer to transform this information into dance motion instructions $\mathcal{T}$, providing initial guidance for motion generation.
+2 and consists of three main components: Dance Instruction Generation, Multi-modal Feature Extraction and Fusion, and Diffusion-based Motion Synthesis.
+The Dance Instruction Generation module (Fig.
+2a) extracts music caption (tempo, beat, key, chord) from the given original music $\mathcal{M}$ and utilizes an LLM as a pseudo-choreographer to transform this information into dance motion instructions $\mathcal{T}$, providing textual guidance for motion generation.
 The Multi-modal Fusion module (Fig.
 2b) then integrates the encoded representations of music $\mathcal{M}$, beats $\mathcal{B}$, and text $\mathcal{T}$ into a unified conditional embedding.
-This fused representation aligns music, beat, and movement intentions, creating a cohesive context for dance generation.
-Finally, the Diffusion-based Motion Synthesis component (Fig.
-2c) uses this unified embedding to guide the generation of realistic, rhythmically aligned dance movements that capture the style and dynamics of the music, producing smooth and expressive dance sequences.
-3.3. Pseudo Instruction Generation
-Figure 3.
-Details of our textual instruction generation process.
-We construct prompts using a templated format.
-The prompt guides an LLM to generate HumanML3D-style dance instructions with constraints on motion focus, perspective, and length.
-The bottom section shows an example of a prompt and its corresponding LLM-generated instruction.
-This module is responsible for transforming music into high-level dance instructions, which serve as semantic guidance for the dance generation process.
-This module comprises two key steps: Music Caption Construction and Textual Instruction Generation.
+This fused representation aligns music, beat, and movement instructions, creating a cohesive context for dance generation.
+Finally, the Diffusion-based Motion Synthesis module (Fig.
+2c) uses this unified embedding to guide the generation of dance movements that capture the style and dynamics of the music.
+3.3 Dance Instruction Generation
+This module is responsible for transforming music into high-level textual dance instructions, which serve as semantic guidance for the dance generation process.
+This module comprises two key steps: Music Caption Construction and LLM-based Instruction Generation.
 Music Caption Construction.
-We extract essential musical attributes from the input music $\mathcal{M}$ to create a comprehensive music caption.
-Following Mustango (Melechovsky et al., 2023), we utilize multiple specialized tools to carry out the extraction:
-  • Tempo: We employ BeatNet (Heydari et al., 2021) to extract tempo information, measured in beats per minute (BPM), to capture the pace of the music.
-This tempo provides a foundational cue for the timing of generated dance movements.
-  • Chord Progression: Using Chordino (Mauch and Dixon, 2010), we identify the sequence of chords within the music.
-Chord progression conveys harmonic changes, which influence the emotional tone of the dance.
-  • Key: The key of the music is extracted using Essentia’s (Bogdanov et al., 2013) KeyExtractor, which determines the tonal center of the piece, providing coherence and structure to the generated dance.
-  • Beat Structure: BeatNet (Heydari et al., 2021) is further utilized to detect beats and downbeats, which serve as rhythmic anchors for synchronizing dance movements with the music.
-Textual Instruction Generation.
-Given the constructed music caption, we generate a textual dance instruction $\mathcal{T}$ that aligns with the corresponding music clip $\mathcal{M}$.
-The instruction serves as a pseudo-guidance for training and is produced using a structured prompt directed to an LLM, which acts as a pseudo choreographer.
-To ensure that the generated instruction is rhythmically and stylistically appropriate, we adopt a carefully designed prompt engineering strategy inspired by the format and motion description style of the HumanML3D dataset (Guo et al., 2022), a comprehensive collection comprising diverse text-motion pair data.
-As shown in Fig 3, we designed a prompt engineering strategy that incorporates several key considerations to guide the model’s output effectively.
-First, we include the duration of the music clip in the prompt, ensuring that generated instructions are aligned with the length and pacing of the music.
-Additionally, to produce genre-specific movement, the prompt specifies a dancing style (e.g., hip-pop, ballet) extracted from data labels, providing the model with style supervision to create dance instructions that match both the music’s rhythm and genre.
-To reduce noise in the generated text, we explicitly direct the model to avoid directional words (e.g., left, right), which can introduce ambiguity or irrelevant details.
-This focus on general body movements without spatial directions helps maintain clarity and relevance in the instructions.
-We also aim to control the text style, structuring the prompt concisely and descriptively.
+We adopt a structured captioning procedure to extract and represent key musical attributes from each input music clip.
+Building upon the framework proposed by Mustango (Melechovsky et al.
+2023), we extract four common symbolic descriptors: tempo, chords, keys, and beat structure using a set of specialized tools.
+To further enhance clarity and generation consistency, we follow a controlled template-based sentence construction process, converting each musical attribute into interpretable textual music captions (e.g., ‘The chord sequence is G.
+The time signature is 4/4.
+The song is played at the pace of Moderato.
+The key of this song is G minor.’).
+These captions are appended to the input prompts to provide rich meta-level guidance for downstream instruction generation.
+LLM-based Instruction Generation.
+To convert music into interpretable guidance for motion, we generate a textual dance instruction $\mathcal{T}$ from the music caption using an instruction-following LLM.
+The LLM acts as a pseudo choreographer, producing stylistically and rhythmically aligned instructions for each music clip.
+These instructions provide explicit descriptions of body movements, offering fine-grained guidance to the motion generator.
+To ensure consistency and controllability, we adopt a carefully designed prompt engineering strategy inspired by the format and motion description style of the HumanML3D dataset (Guo et al.
+2022).
+As shown in Fig 3, we incorporate key elements such as clip duration and dance style.
+We also avoid directional words and overly fine-grained spatial cues to improve clarity and reduce ambiguity.
+Figure 3: Details of our LLM-based textual instruction generation process. The prompt guides an LLM to generate HumanML3D-style dance instructions.
 Instruction Results.
 The output of this process is a set of natural language dance instructions that provide semantic cues for the dance generation model.
-For instance, given a music track with a moderate tempo and a hip-hop style, the generated instruction might read: The dancer bounces lightly to the beat, alternating shoulder pops with a subtle chest pump.
-These instructions capture both the rhythmic alignment and stylistic characteristics of the music, bridging the semantic gap between audio and motion.
-Example is shown in Fig.
-3.
-3.4. Multi-modal Feature Extraction and Fusion
-This module integrates multiple conditions from text, music, and beat information into a unified conditional embedding.
-We first encode each modality individually, then hierarchically fuse these embeddings to leverage their complementary strengths effectively.
-3.4.1. Multi-modal Feature Extraction
-Music Encoding.
-We leverage a pre-trained Jukebox (Dhariwal et al., 2020) model to extract high-level musical features from raw audio inputs.
-Specifically, we input the raw audio into the frozen Jukebox encoder, obtaining embeddings that represent the musical content.
-To adapt these embeddings for our task, we introduce a trainable two-layer Transformer encoder that processes the Jukebox-derived features, allowing the model to learn task-specific representations while retaining the rich musical information captured by Jukebox, resulting in embeddings $E_{M}\in\mathbb{R}^{d_{M}}$.
-Text Encoding.
-The dance instruction generated by LLMs is processed through a CLIP (Radford et al., 2021) text encoder to capture semantic details about body movements and style.
-To tailor the text embeddings to our specific application, we incorporate a learnable 4-layer Transformer as a conditioning encoder.
-This additional Transformer processes the frozen CLIP embeddings, enabling the model to capture extra semantic dance guidance.
-To better leverage prior knowledge from existing text-to-motion models, we initialize our entire text encoder with the pretrained weights from MotionDiffuse (Zhang et al., 2022).
-This is motivated by the fact that our pseudo instructions are generated to closely follow the style and vocabulary of the HumanML3D (Guo et al., 2022) dataset, which was also used to train MotionDiffuse.
-Since our textual instructions are automatically generated and no paired text-motion supervision is available, we do not apply direct supervised training to the text encoder.
-Instead, we rely on the strong prior learned from MotionDiffuse to provide a semantically meaningful representation.
-The pretrained CLIP and Transformer modules together form our text encoder, and the final textual embedding is denoted as $E_{T}\in\mathbb{R}^{d_{T}}$.
-Beat Encoding.
-In addition to basic musical features, we explicitly incorporate beat proximity information to enhance rhythmic awareness in generated motion.
-We apply Librosa (McFee et al., 2015) to extract both beat and downbeat positions from the audio.
-These positions are discrete time points that mark the perceived rhythmic structure of the music.
-To represent rhythmic proximity in a continuous form, we follow a Gaussian-based approach: for each frame, we compute a Gaussian-like pulse centered around the nearest beat and downbeat positions.
-This encoding reflects the temporal distance to nearby rhythmic accents, enabling the model to perceive and respond to the rhythmic context over time.
-We then project these proximity values into a learnable embedding space using a linear projection layer, resulting in beat embeddings $E_{B}\in\mathbb{R}^{d_{B}}$.
-Together, these extracted features capture the unique attributes of each modality, providing a foundation for effective cross-modal integration.
-Each of these extracted feature embeddings provides distinct complementary information: textual embedding $E_{T}$ offers semantic pseudo guidance, music embedding $E_{M}$ captures musical style, and beat embedding $E_{B}$ encodes rhythmic timing.
-3.4.2. Multi-modal Hierarchical Fusion
-We fuse these modality embeddings through a hierarchical, two-level strategy designed to incrementally integrate rhythmic, musical, and textual guidance.
-Musical-Rhythmic Fusion. Initially, we fuse music embeddings $E_{M}$ with rhythmic beat embedding $E_{B}$ using an additive fusion approach:
-(1) $E_{M}^{\prime}=E_{M}+E_{B},$
-This straightforward yet effective fusion explicitly integrates rhythmic timing information directly into the music representations.
-By enhancing music embeddings with beat cues, it enables the model to accurately capture rhythmically significant moments within the music.
-Music-Text Semantic Fusion.
-Subsequently, we incorporate textual information from dance instruction embeddings $E_{T}$ to provide semantic guidance for dance generation.
-Specifically, we concatenate the textual embeddings $E_{T}$ with the previously fused musical-rhythmic embeddings $E_{M}^{\prime}$:
-(2) $E_{F}=[E_{M}^{\prime};E_{T}],$
-generating the final unified conditional embedding $E_{F}$.
-This second fusion stage integrates high-level choreographic instructions, derived from natural language guidance, with detailed rhythmic-musical context.
-Thus, the final embedding $E_{F}$ effectively combines semantic choreographic concepts from textual instructions and detailed rhythmic structure from music and beat embeddings.
-The final conditional embedding $E_{F}$ encapsulates a balanced representation of the text, music, and beat, ensuring that generated dance movements are not only synchronized with the rhythm but also stylistically aligned and semantically meaningful.
-This fused embedding serves as the input for the diffusion-based motion synthesis module, enabling the generation of realistic, expressive dance sequences that reflect the musical nuances and choreographic intentions embedded in the text.
-3.5. Diffusion-based Motion Synthesis
-This module leverages a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) to progressively guide noisy samples toward realistic dance motions, conditioned on the multi-modal embedding $E_{F}$ from the previous Multi-modal Fusion stage.
-We model dance generation as a Markovian diffusion process, in which a sequence of noisy latent representations is gradually denoised to reveal the desired motion.
-The forward diffusion process begins with a clean motion sequence $\mathcal{X}$ and iteratively adds Gaussian noise over $T$ timesteps, eventually transforming it into a nearly random noise sample $z_{T}$.
-Formally, at each timestep $t$, the forward process is defined as:
-(3) $\mathit{q(z_{t}|z_{t-1})}=\mathcal{N}(\sqrt{\mathit{1-\beta_{t}}}\mathit{z_{t-1}},\beta_{t}\mathit{I}),$
-where $\beta_{t}\in(0,1)$ is a variance schedule parameter that controls the amount of noise added at each step.
-By the final timestep $T$, the noisy sequence $z_{T}$ closely approximates a standard Gaussian distribution $\mathcal{N}(0,\mathit{I})$.
-To recover the original motion from this noisy representation, we employ a reverse diffusion process.
-This reverse process begins with the noisy sample $z_{T}$ and iteratively denoises it, guided by the conditional embedding $E_{F}$, to approximate the original sequence $\mathcal{X}$.
-At each step $t$, we predict the previous timestep $z_{t-1}$ from $z_{t}$ using:
-(4) $\mathit{p_{\theta}(z_{t-1}|z_{t},E_{F})}=\mathcal{N}(\mathit{z_{t-1}};\mu_{\theta}(z_{t},t,E_{F}),\sigma_{t}^{2}\mathit{I}),$
-where $\mu_{\theta}(z_{t},t,E_{F})$ represents the predicted mean for $z_{t-1}$ based on the model parameters $\theta$, and $\sigma_{t}$ is a learned variance that adjusts the noise schedule during the denoising process.
-We adopt the simplified training objective from DDPM (Ho et al., 2020), minimizing the mean squared error (MSE) between the model’s predictions and the true motion data across all timesteps.
-This objective encourages the model to reconstruct realistic dance sequences from noisy inputs while aligning with the semantic and rhythmic cues encoded in $E_{F}$.
-Formally, the training objective is defined as:
-(5) $\mathcal{L}_{\mathrm{simple}}=\mathbb{E}_{\mathbf{x},t}[\parallel\mathbf{x}-\mathbf{\hat{x}}(z_{t},t,E_{F})\parallel^{2}_{2}],$
-where $\mathbf{\hat{x}}(z_{t},t,E_{F})$ denotes the reconstructed motion at timestep $t$ conditioned on the multi-modal embedding $E_{F}$.
-3.6. Loss Functions
+For instance (Fig.
+3), given a music track with a moderate tempo and a Hip-Hop style, the generated instruction might read: This person bounces with the knees while alternating shoulder pops and chest pops, adding a quick head nod on every beat.
+These instructions capture both the rhythmic alignment and stylistic characteristics of the music, bridging the semantic gap between music and motion.
+3.4 Multi-modal Feature Extraction and Fusion
 Method PFC $\downarrow$ $\mathrm{Div}_{k}\rightarrow$ $\mathrm{Div}_{g}\rightarrow$ Beat Align. Score $\uparrow$
 Ground Truth 1.332 10.61 7.48 0.24
-FACT (Li et al., 2021) 2.254 10.85 6.14 0.22
-Bailando (Siyao et al., 2022) 1.754 7.92 7.72 0.23
-EDGE (Tseng et al., 2023) 1.536 9.48 5.72 0.23
-BADM (Zhang et al., 2024b) 1.424 8.29 6.76 0.24
-Beat-it (Huang et al., 2024) 0.966 9.66 7.25 0.66
+FACT (Li et al. 2021) 2.254 10.85 6.14 0.22
+Bailando (Siyao et al. 2022) 1.754 7.92 7.72 0.23
+EDGE (Tseng, Castellon, and Liu 2023) 1.536 9.48 5.72 0.23
+BADM (Zhang et al. 2024a) 1.424 8.29 6.76 0.24
+Beat-it (Huang et al. 2024) 0.966 9.66 7.25 0.66
 Ours DanceChat 0.828 10.64 7.36 0.27
-Table 1.
-Quantitative comparison with state-of-the-art methods on the AIST++ dataset.
+Table 1: Quantitative comparison with state-of-the-art methods on the AIST++ dataset.
 FACT and Bailando are Transformer- and VQ-VAE-based models that require a starting pose.
@@
 Our method outperforms the other methods on $\mathrm{PFC}$, $\mathrm{DIV}_{k}$ and $\mathrm{DIV}_{g}$, and performs the second-best on $\mathrm{BAS}$.
+This module integrates multiple conditions from text, music, and beat information into a unified conditional embedding.
+We first encode each modality individually, then hierarchically fuse these embeddings to leverage their complementary strengths effectively.
+Multi-modal Feature Extraction
+Music Encoding.
+We leverage a pre-trained Jukebox (Dhariwal et al.
+2020) model to extract high-level musical features from raw audio inputs.
+Specifically, we input the raw audio into the frozen Jukebox encoder, obtaining embeddings that represent the musical content.
+To adapt these embeddings for our task, we introduce a trainable two-layer Transformer encoder that processes the Jukebox-derived features, allowing the model to learn task-specific representations while retaining the rich musical information captured by Jukebox, resulting in embeddings $E_{M}\in\mathbb{R}^{d_{M}}$.
+Text Encoding.
+We process the LLM-generated dance instructions using a CLIP-based text encoder (Radford et al.
+2021) followed by a trainable Transformer to extract task-specific embeddings.
+To enhance semantic alignment, we initialize the encoder with pretrained weights from MotionDiffuse (Zhang et al.
+2024b), which is trained on HumanML3D-style (Guo et al.
+2022) descriptions.
+This setup enables our model to interpret dance instructions effectively, without requiring paired text-motion supervision.
+The final textual embedding is denoted as $E_{T}\in\mathbb{R}^{d_{T}}$.
+Beat Encoding.
+To enhance rhythmic awareness in generated motion, we explicitly incorporate beat proximity features.
+We use Librosa (McFee et al.
+2015) to extract beat and downbeat positions from audio.
+These discrete time points mark the perceived rhythmic structure of music.
+To convert this into a continuous representation, we compute a Gaussian-shaped pulse for each frame centered around the nearest beat/downbeat.
+This reflects the frame’s temporal distance to nearby rhythmic accents, enabling the model to perceive and respond to rhythm over time.
+We then project these proximity values into a learnable embedding space using a linear projection layer, resulting in beat embeddings $E_{B}\in\mathbb{R}^{d_{B}}$.
+Together, the extracted features from all modalities form a unified representation: the textual embedding $E_{T}$ provides pseudo motion guidance, the musical embedding $E_{M}$ captures style and structure, and the beat embedding $E_{B}$ encodes rhythmic timing, enabling effective and complementary cross-modal integration.
+Multi-modal Hierarchical Fusion
+We fuse these modality embeddings through a hierarchical, two-level strategy designed to incrementally integrate rhythmic, musical, and textual guidance.
+Music-Beat Rhythmic Fusion. Initially, we fuse music embeddings $E_{M}$ with rhythmic beat embedding $E_{B}$ using an additive fusion approach:
+$E_{M}^{\prime}=E_{M}+E_{B},$ (1)
+This straightforward yet effective fusion explicitly integrates rhythmic timing information directly into the music representations.
+By enhancing music embeddings with beat cues, it enables the model to accurately capture rhythmically significant moments within the music.
+Music-Text Semantic Fusion.
+Subsequently, we incorporate textual information from dance instruction embeddings $E_{T}$ to provide semantic guidance for dance generation.
+Specifically, we concatenate the textual embeddings $E_{T}$ with the previously fused musical-beat embeddings $E_{M}^{\prime}$:
+$E_{F}=[E_{M}^{\prime};E_{T}],$ (2)
+generating the final unified conditional embedding $E_{F}$.
+This second fusion stage integrates high-level choreographic instructions, derived from natural language guidance, with detailed rhythmic-musical context.
+Thus, the final embedding $E_{F}$ effectively combines semantic choreographic concepts from textual instructions and detailed rhythmic structure from music and beat embeddings.
+The unified embedding $E_{F}$ is then used to condition the diffusion-based motion synthesis module, guiding the generation of realistic and expressive dance that reflects both musical structure and textual intent.
+3.5 Diffusion-based Motion Synthesis
+Diffusion Model
+This module leverages a denoising diffusion probabilistic model (DDPM) (Ho, Jain, and Abbeel 2020) to progressively guide noisy samples toward realistic dance motions, conditioned on the multi-modal embedding $E_{F}$ from the previous Multi-modal Fusion stage.
+We model dance generation as a Markovian diffusion process.
+The forward diffusion process begins with a clean motion sequence $\mathbf{x}_{0}$ and iteratively adds Gaussian noise over $T$ timesteps, eventually transforming it into a nearly random noise sample $\mathbf{z}_{T}$.
+Formally, at each timestep $t$, the forward process is defined as:
+$\mathit{q}(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\sqrt{\mathit{1-\beta_{t}}}\mathbf{z}_{t-1},\beta_{t}\mathit{I}),$ (3)
+where $\beta_{t}\in(0,1)$ is a variance schedule parameter that controls the amount of noise added at each step.
+By the final timestep $T$, the noisy sequence $\mathbf{z}_{T}$ closely approximates a standard Gaussian distribution $\mathcal{N}(0,\mathit{I})$.
+To recover the original motion from this noisy representation, we employ a reverse diffusion process.
+This reverse process begins with the noisy sample $\mathbf{z}_{T}$ and iteratively denoises it, guided by the conditional embedding $E_{F}$, to approximate the original sequence $\mathbf{x}_{0}$.
+We adopt the simplified training objective from DDPM (Ho, Jain, and Abbeel 2020), minimizing the mean squared error (MSE) between the model’s predictions and the true motion data across all timesteps.
+Formally, the training objective is defined as:
+$\mathcal{L}_{\mathrm{simple}}=\mathbb{E}_{\mathbf{x},t}[\parallel\mathbf{x}_{0}-\mathbf{\hat{x}}_{\theta}(\mathbf{z}_{t},t,E_{F})\parallel^{2}_{2}],$ (4)
+where $\mathbf{\hat{x}}_{\theta}(\mathbf{z}_{t},t,E_{F})$ denotes the reconstructed motion at timestep $t$ conditioned on the multi-modal embedding $E_{F}$ with model parameters $\theta$.
+Loss Functions
 Kinematic Loss.
 To enhance the physical plausibility of generated motions, we incorporate several auxiliary losses commonly used in kinematic motion generation.
-In addition to the basic reconstruction loss $\mathcal{L}_{\mathrm{simple}}$, we adopt auxiliary terms introduced by Tevet et al.(Tevet et al., 2022b) and further refined in EDGE(Tseng et al., 2023), targeting three key aspects of motion realism: joint positions (Eq.
-6), joint velocities (Eq.
-7), and foot-ground contact consistency (Eq.
-8).
+In addition to the basic reconstruction loss $\mathcal{L}_{\mathrm{simple}}$, we adopt auxiliary terms introduced by Tevet et al.(Tevet et al.
+2022) and further refined in EDGE(Tseng, Castellon, and Liu 2023), targeting three key aspects of motion realism: joint positions (Eq.
+5), joint velocities (Eq.
+6), and foot-ground contact consistency (Eq.
+7).
 Below, $\mathbf{FK(\cdot)}$ denotes the forward kinematic function that converts joint angles into joint positions and the $\mathit{i}$-th superscript denotes the frame index.
@@
 7 only applies to the foot joints, and $\mathbf{\hat{b}}_{i}$ is binary foot contact signal.
-(6) $\mathcal{L}_{\mathrm{joint}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}\parallel\mathbf{FK}(\mathbf{x}_{i})-\mathbf{FK}(\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$
-(7) $\mathcal{L}_{\mathrm{vel}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{x}_{i+1}-\mathbf{x}_{i})-(\mathbf{\hat{x}}_{i+1}-\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$
-(8) $\mathcal{L}_{\mathrm{contact}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{FK}(\mathbf{x}_{i+1})-\mathbf{FK}(\mathbf{\hat{x}}_{i}))\cdot\mathbf{\hat{b}}_{i}\parallel^{2}_{2}$
+$\mathcal{L}_{\mathrm{joint}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}\parallel\mathbf{FK}(\mathbf{x}_{i})-\mathbf{FK}(\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$ (5)
+$\mathcal{L}_{\mathrm{vel}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{x}_{i+1}-\mathbf{x}_{i})-(\mathbf{\hat{x}}_{i+1}-\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$ (6)
+$\mathcal{L}_{\mathrm{contact}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{FK}(\mathbf{x}_{i+1})-\mathbf{FK}(\mathbf{\hat{x}}_{i}))\cdot\mathbf{\hat{b}}_{i}\parallel^{2}_{2}$ (7)
 Multi-modal Alignment Loss.
-To bridge the semantic gap between music and motion, we design a multi-modal alignment loss function that leverages text as an intermediate modality.
-Unlike traditional approaches that directly align music and motion, our method calculates similarity between motion-text and text-music pairs at each timestep $\mathit{t}$ in the diffusion process.
-By using text as a common semantic space, we allow the model to capture descriptive and interpretable features that relate both to the audio’s rhythmic and tonal characteristics and to the motion’s dynamic and stylistic attributes.
-Specifically, for each timestep, we compute cosine similarity between motion and text embeddings and text and music embeddings, promoting alignment across all three modalities through the text-based intermediary.
-The total alignment loss is defined as Eq.
-9, encouraging consistent semantic alignment between music and motion via the shared text representation.
-(9) $\mathcal{L}_{\mathrm{align}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}[(1-\mathrm{sim}(E_{t_{i}},E_{m_{i}}))+(1-\mathrm{sim}(E_{t_{i}},E_{x_{i}}))]$
-This alignment loss is then combined with the other auxiliary losses, weighted by factors $\lambda$ to balance the magnitudes of the losses at the start of training.
-(10) $\begin{split}\mathcal{L}&=\mathcal{L}_{\mathrm{simple}}+\lambda_{\mathrm{joint}}\mathcal{L}_{\mathrm{joint}}+\lambda_{\mathrm{vel}}\mathcal{L}_{\mathrm{vel}}+\\ &\lambda_{\mathrm{contact}}\mathcal{L}_{\mathrm{contact}}+\lambda_{\mathrm{align}}\mathcal{L}_{\mathrm{align}}\end{split}$
-4. Experiments
-4.1. Experiment Setup
-Figure 4.
-Example results generated by our model.
+To bridge the semantic gap between music and motion, we design a multi-modal alignment loss that leverages text as an intermediate modality.
+Unlike prior methods that directly align music and motion, we instead calculates similarity between motion-text and text-music pairs at each timestep $\mathit{t}$ in the diffusion process.
+By using text as a shared semantic space, the model captures interpretable features related to both musical cues and motion dynamics.
+Specifically, for each timestep, we compute cosine similarity between motion-text embeddings and text-music embeddings, promoting alignment across all three modalities through the text-based intermediary.
+The total alignment loss is defined as follow:
+$\mathcal{L}_{\mathrm{align}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}[(1-\mathrm{sim}(E_{t_{i}},E_{m_{i}}))+(1-\mathrm{sim}(E_{t_{i}},E_{x_{i}}))]$ (8)
+This loss encourages consistent semantic alignment between music and motion via text.
+It is combined with the other auxiliary losses, weighted by factors $\lambda$ to balance the magnitudes of the losses during training.
+$\begin{split}\mathcal{L}&=\mathcal{L}_{\mathrm{simple}}+\lambda_{\mathrm{joint}}\mathcal{L}_{\mathrm{joint}}+\lambda_{\mathrm{vel}}\mathcal{L}_{\mathrm{vel}}+\\ &\lambda_{\mathrm{contact}}\mathcal{L}_{\mathrm{contact}}+\lambda_{\mathrm{align}}\mathcal{L}_{\mathrm{align}}\end{split}$ (9)
+4 Experiments
+4.1 Experiment Setup
+Figure 4: Example results generated by our model.
 Each sequence illustrates a 3D dance motion generated from a given music clip, conditioned on an LLM-generated textual instruction.
@@
 Dataset.
-We use the AIST++(Li et al., 2021) dataset for both training and testing.
-This dataset comprises 1,408 dance sequences paired with music, covering 10 distinct street dance genres.
-Each genre is synchronized with music from 6 different tracks, each with unique beats per minute (BPM) and musical characteristics, allowing for a variety of rhythmic and stylistic alignment in the generated dance motions.
-All dance motions are represented as 3D poses captured at 60 FPS, providing detailed joint-level information for each frame.
-For our experiments, we segment these music-motion pair sequences into 5-second clips sampled at 30 FPS, and extract music caption and textual guidance based on each segmented music clip.
-We follow the standard train/test splits provided in the original dataset.
+We use the AIST++ (Li et al.
+2021) dataset for both training and testing, which contains 1,408 music-dance pairs across 10 street dance genres.
+Each genre is paired with 6 distinct music tracks, enabling rhythmic and stylistic variation.
+We follow the standard train/test split and segment the sequences into 5-second clips at 30 FPS.
+Music captions and textual guidance are generated based on each segmented clip.
 Implementation details.
-In the pseudo instruction generation module, we utilize the OpenAI ChatGPT-4o API (OpenAI API, [n.
-d.]) to generate textual dance motion instruction based on pre-extracted music captions.
-For prompt engineering, we use the annotated genre information during training.
-At test time, we replace the genre placeholder in the prompt with the literal word any.
-In the multi-modal feature extraction and fusion module, we adopt an encoding approach in which pre-trained models are frozen to preserve general representations, while lightweight trainable layers are added for task-specific adaptation.
-We first extract features using a frozen Jukebox model and subsequently apply two additional Transformer encoder layers to refine the temporal and semantic representations.
-For text encoding, we use the frozen text encoder from CLIP ViT-B/32 (Radford et al., 2021) followed by another two-layer Transformer encoder to adapt the output for conditional guidance.
+In the dance instruction generation module, we utilize the ChatGPT-4o API (OpenAI API ) to generate textual dance motion instruction based on pre-extracted music captions.
+For music encoding, we extract features using a frozen Jukebox (Dhariwal et al.
+2020) model and subsequently apply two additional Transformer encoder layers to refine the temporal and semantic representations.
+For text encoding, we use the frozen text encoder from CLIP ViT-B/32 (Radford et al.
+2021) followed by another 4-layer Transformer encoder to adapt the output for conditional guidance.
 The latent dimensionality of both the music and text encoders is set to 512.
 Our model has 54M total parameters, and was trained on 1 NVIDIA A100 GPU for 48 hours with a batch size of 128.
-4.2. Comparison to Existing Methods
-We have selected FACT (Li et al., 2021), Bailando (Siyao et al., 2022), EDGE (Tseng et al., 2023), BADM (Zhang et al., 2024b) and Beat-it (Huang et al., 2024) for comparisons with our DanceChat method.
-FACT utilizes a full-attention mechanism to capture dependencies across music and motion, leveraging autoregressive generation to maintain consistency in dance sequences.
-Bailando builds upon this by incorporating a VQ-VAE (Van Den Oord et al., 2017) and GPT (Radford, 2018) framework, enhanced with an actor-critic reinforcement learning scheme to improve dance quality.
-EDGE introduces a transformer-based diffusion model, serving as the backbone of our method.
-BADM, as a follow-up to EDGE, proposes a bidirectional autoregressive diffusion model that enforces harmony in both forward and backward directions, capturing local motion details for smoother dance sequences.
-Similarly, Beat-it extends EDGE by presenting a beat-synchronized, key pose-guided 3D dance generation framework, integrating explicit beat awareness and key pose guidance to ensure alignment between dance motions and musical beats.
-To evaluate the effectiveness of our approach, we conduct a series of experiments to assess the performance in terms of physical plausibility, motion diversity, and beat alignment score (BAS), which together evaluate the physical realism, stylistic variability, and rhythmic quality of the generated dance motion.
-The quantitative metrics are shown in Table 1.
-In addition to this, we also conduct a human evaluation to gather subjective feedback on the overall quality and coherence of the generated motions, as shown in Table 2.
+4.2 Comparison to Existing Methods
+Method Ours (AIST++) Ours (In-the-Wild)
+FACT 86.25% N/A
+Bailando 61.88% N/A
+EDGE 60.00% 66.36%
+Ground Truth 48.12% N/A
+Table 2: User study results on AIST++ test set and in-the-wild music.
+We present the Preference Rate $\uparrow$ of user evaluations that prefer DanceChat (Ours) to FACT (Li et al.
+2021), Bailando (Siyao et al.
+2022), and EDGE (Tseng, Castellon, and Liu 2023).
+We have selected FACT (Li et al.
+2021), Bailando (Siyao et al.
+2022), EDGE (Tseng, Castellon, and Liu 2023), BADM (Zhang et al.
+2024a) and Beat-it (Huang et al.
+2024) for comparisons with our DanceChat method.
+FACT applies a full-attention mechanism for cross-modal dependency modeling, using autoregressive generation for temporal consistency.
+Bailando enhances this with a VQ-VAE (Van Den Oord, Vinyals et al.
+2017) and GPT (Radford 2018)-based framework, together with actor-critic reinforcement learning.
+EDGE adopts a transformer-based diffusion model, which serves as the backbone of our method.
+BADM builds on EDGE by proposing a bidirectional autoregressive diffusion model for smoother local motion.
+Beat-it further extends EDGE by integrating beat supervision and key-pose guidance to improve rhythm-movement alignment.
+To evaluate the effectiveness of our approach, we conduct a series of experiments to assess the performance in terms of physical plausibility, motion diversity, and beat synchronization.
+The quantitative metrics are shown in Tab.
+1.
+In addition to this, we also conduct a user study to gather subjective feedback on the overall quality and coherence of the generated motions, as shown in Tab.
+2.
 Motion Quality.
-We incorporate the Physical Foot Contact score (PFC) (Tseng et al., 2023) to evaluate the foot-ground interactions, a critical component of physical realism in dance generation.
-The PFC score assesses the accuracy of foot contact behavior, ensuring that generated sequences avoid unrealistic foot sliding and maintain natural center of mass (COM) dynamics.
-A lower PFC score indicates that the generated movements more closely approximate realistic foot-ground interactions, with stable foot placements and controlled COM dynamics, leading to improved balance and physical plausibility.
-As shown in Table 1, DanceChat achieves the best performance in $\mathrm{PFC}$, with an improvement of 0.138 compared with SOTA method Beat-it (Huang et al., 2024).
-This result highlights DanceChat’s ability to generate dance sequences with high physically realistic foot-ground interactions.
+We use the Physical Foot Contact (PFC) score (Tseng, Castellon, and Liu 2023) to evaluate foot-ground interactions, a key factor for physical realism.
+It measures the accuracy of foot contact, ensuring that generated sequences avoid unrealistic foot sliding and maintain natural center of mass (COM) dynamics.
+Lower scores indicate more stable and realistic foot-ground interactions.
+As shown in Tab.
+1, DanceChat achieves the best PFC performance, improving by 0.138 over SOTA method Beat-it.
+This highlights our model’s ability to produce physically plausible dance sequences.
 Motion Diversity.
-To assess the diversity of the generated dance motions, we follow the approach outlined by Bailando (Siyao et al., 2022).
-We compute the mean Euclidean distance within the motion feature space to quantify the spread of generated motions.
-Specifically, we calculate two diversity metrics: $\mathrm{Div}_{k}$, which reflects diversity in the kinematic feature space, and $\mathrm{Div}_{g}$, which measures diversity in the geometric feature space.
-These metrics are computed over 5-second dance clips to maintain consistency with prior benchmarks (Tseng et al., 2023; Zhang et al., 2024b; Huang et al., 2024).
-Following observations from EDGE (Tseng et al., 2023), excessively jittery or unstable motions can lead to artificially inflated diversity scores, occasionally exceeding those of the ground truth.
-Therefore, rather than maximizing diversity, models should aim to match the distributional spread of the ground truth, indicating an appropriate balance between variation and realism.
-As shown in Table 1, our approach achieves a $\mathrm{Div}_{g}$ score of 7.36 and a $\mathrm{Div}_{k}$ score of 10.64, which are closer to the ground truth than all the other methods.
-We attribute to our model’s use of global diffusion techniques and fundamental dance primitives, enabling it to learn diverse choreography patterns.
+To assess diversity, we compute the mean Euclidean distance of motion features across 5-second clips, following standard metrics $\mathrm{Div}_{k}$ (kinematic space) and $\mathrm{Div}_{g}$ (geometric space) outlined by Bailando.
+Our metrics are consistent with prior work and computed to match the distributional spread of the ground truth, avoiding the inflation from jittery or unstable motion.
+As shown in Tab.
+1, our approach achieves a $\mathrm{Div}_{g}$ of 7.36 and $\mathrm{Div}_{k}$ of 10.64, both closest to ground truth among all methods, reflecting a good balance between variation and realism.
 Beat Alignment Score (BAS).
-To evaluate the rhythmic alignment of generated dance movements with the accompanying music, we use the BAS metric, which measures the consistency between the music beats and the kinematic beats of the generated motion.
-Following Bailando (Siyao et al., 2022), this metric rewards alignment when local minima in joint speed (indicative of pauses or transitions in motion) correspond with the musical beats.
-As shown in Table 1, our model achieves the second-best BAS score, indicating a strong ability to follow the beat and maintain rhythmic coherence.
-The highest BAS is achieved by Beat-it (Huang et al., 2024), which explicitly optimizes for beat alignment using a strong beat-supervised loss combined with key-pose constraints, allowing it to tightly synchronize motions with musical beats.
-In contrast, our model achieves competitive rhythmic alignment without requiring strong beat-level supervision.
-Instead, we rely on a flexible multi-modal fusion framework that incorporates both musical and beat features in a weakly supervised manner.
-This enables our model to produce rhythmically aware motion while maintaining better generalization across genres and motion styles.
-Human Evaluation.
-To gain a deeper understanding of the visual and rhythmic quality of our method, we conduct a user study comparing the dance sequences generated by DanceChat with those produced by baseline methods on the AIST++ dataset and the in-the-wild musics.
-This study involves 30 participants, each engaging in individual evaluations.
-Participants are shown a series of 20 pairs of comparison videos, each approximately 10 seconds in length.
-Each pair includes one video generated by DanceChat and one by a competing method, synchronized to the same music track.
-The participants’ task is to identify which of the two videos feel better.
-The results of the user study on AIST++ are summarized in Table 2.
-Notably, DanceChat outperformed all the baseline results.
-We also conduct a separate evaluation in an “in-the-wild” setting, using unseen music tracks to test DanceChat’s generalization capabilities.
-As shown in Table 3, DanceChat achieves a preference rate of 66.36% over EDGE in these real-world scenarios, further confirming its robustness and adaptability across different musical contexts.
-In addition to the user study, an observable reduction in foot sliding is noted in the DanceChat videos, further confirming the PFC improvement.
-4.3. Ablation Study
-We perform an ablation study on the AIST++ dataset to evaluate the impact of different modalities (music, beat, and text) and the effectiveness of the multi-modal alignment loss in generating high-quality dance motions.
-The results, shown in Table 4, are reported in terms of $\mathrm{PFC}$ (for motion realism), $\mathrm{DIV}_{k}$ (for motion diversity), and BAS (for beat alignment score), with lower $\mathrm{PFC}$ and higher $\mathrm{DIV}_{k}$ and BAS indicating better performance.
+We use the BAS metric to evaluate the rhythmic alignment of generated dance with music, which measures the consistency between musical beats and joint speed minima in motion.
+As shown in Tab.
+1, our model achieves the second-best BAS score, demonstrating strong beat-tracking ability.
+While Beat-it, which uses beat-supervised training, achieves the highest score, our method obtains competitive performance without explicit beat-level supervision.
+This is enabled by our flexible multi-modal fusion framework, which jointly encodes music and beat features to produce rhythm-aware motion and generalize across styles.
+User Study.
+To assess the visual and rhythmic quality of our method, we conduct a user study comparing DanceChat with baseline methods on AIST++ and in-the-wild music.
+30 participants each evaluated 20 pairs of 10s videos, where each pair included a DanceChat result and one from a competing method, synchronized to the same music.
+Participants were asked to select the more appealing video.
+Results are shown in Tab.
+2.
+DanceChat outperforms all baselines on AIST++.
+In the in-the-wild setting, it achieves a 66.36% preference over EDGE, confirming its robustness to unseen music.
+Additionally, an observable reduction in foot sliding is noted, further confirming the PFC improvement.
+4.3 Ablation Study
+We perform an ablation study on the AIST++ to evaluate the impact of different modalities and the effectiveness of our multi-modal alignment loss.
+Results in Tab.
+3 are reported using $\mathrm{PFC}$, $\mathrm{DIV}_{k}$, and BAS, where lower $\mathrm{PFC}$ and higher others indicate better performance.
 Impact of Modality Inputs.
 To understand the individual contributions of each modality, we assess various configurations by selectively including or excluding the music (M), beat (B), and text (T) modalities.
-The baseline model, using only music (M), performs the lowest across all metrics, with limited diversity and weak rhythmic alignment.
+The baseline model, using only music (M), performs the worst, showing limited diversity and weak rhythmic alignment.
 Adding beat information (M+B) enhances rhythmic alignment, indicating that beat cues contribute both to physical realism in generated movement patterns and to synchronization with the music.
@@
 However, lacking beat information, this configuration underperforms in BAS.
-Interestingly, combining all three modalities (M+B+T) produces the highest BAS and consistently strong results in other metrics.
+Combining all three modalities (M+B+T) produces the highest BAS and consistently strong results in other metrics.
 Impact of Multi-Modal Alignment Loss.
@@
 Loss), which leverages text as an intermediary bridge between music and motion to narrow the semantic gap between these modalities.
-The alignment loss leads to both improved $\mathrm{PFC}$) and $\mathrm{DIV}_{k}$, demonstrating its effectiveness in facilitating cross-modal integration.
+The alignment loss improves $\mathrm{PFC}$) and $\mathrm{DIV}_{k}$, demonstrating its effectiveness in cross-modal integration.
 This suggests that using text as a bridge in the alignment loss helps create semantically rich and rhythmically aligned dance motions that benefit from the strengths of each modality without compromising diversity or quality.
-[…]
-FACT (Li et al., 2021) 86.25%
-Bailando (Siyao et al., 2022) 61.88%
-EDGE (Tseng et al., 2023) 60.00%
-Ground Truth 48.12%
-Table 2.
-User study results on AIST++ test set.
-We present the percentage of user evaluations that prefer DanceChat to FACT (Li et al., 2021), Bailando (Siyao et al., 2022), and EDGE (Tseng et al., 2023).
-[…]
-EDGE (Tseng et al., 2023) 66.36%
-Table 3. User study results of in-the-wild music. We compare DanceChat against EDGE(Tseng et al., 2023).
 Ablations Metrics
-M B T PFC $\downarrow$ Divk $\rightarrow$ Beat Align. Score $\uparrow$
-1.536 9.48 0.26
+M B T PFC $\downarrow$ Divk $\rightarrow$ BAS $\uparrow$
+1.536 9.48 0.23
 1.066 8.76 0.27
@@
 Ground Truth 1.332 10.61 0.24
-Table 4. Ablation study of modalities, where M indicates music, B indicates beat and T indicates text. The best results are indicated in bold, and the second best are underlined.
-5. Conclusion
-In this paper, we present DanceChat, a multi-modal, language-guided framework for music-to-dance generation, addressing the significant semantic gap between music and dance.
-By leveraging a Large Language Model (LLM) as a pseudo-choreographer, our approach introduces natural language as an intermediary between music and motion, transforming abstract musical elements into concrete dance instructions.
-Our proposed model architecture comprises three core components: Pseudo Instruction Generation, which translates musical meta-information into text-based dance instructions; Multi-modal Fusion, which integrates music, beat, and textual guidance into a cohesive representation; and Diffusion-based Motion Synthesis, which employs a denoising process to generate smooth, expressive dance motions.
-Through comprehensive experiments on the AIST++ dataset, DanceChat demonstrates competitive performance across multiple metrics.
-The results confirm that each modality contributes uniquely to dance quality, while our multi-modal alignment loss plays a crucial role in harmonizing these modalities to produce musically aligned sequences.
-Limitations and Future Work.
-While DanceChat achieves strong results, certain limitations remain.
-First, the reliance on pre-trained LLMs and external encoders introduces a dependency on existing models that may not be fully optimized for dance generation.
-Future work could explore custom training for language models specifically tailored for dance, potentially enhancing the semantic richness of generated instructions.
+Table 3: Ablation study of modalities, where M indicates music, B indicates beat and T indicates text. The best results are indicated in bold, and the second best are underlined.
+5 Conclusion
+We present DanceChat, a novel multi-modal framework that leverages language as an intermediate modality to bridge the semantic gap between music and dance.
+By incorporating textual dance instructions generated by LLMs and integrating them with both music and beat features, our approach enables semantically guided and rhythmically aligned motion generation.
+Comprehensive experiments on AIST++ dataset demonstrate that DanceChat achieves strong performance across multiple aspects of dance quality, including physical plausibility, motion diversity, and beat synchronization.
+Ablation results confirm that each modality contributes uniquely to dance quality, while our multi-modal alignment loss plays a crucial role in harmonizing these modalities to produce musically aligned motion sequences.
+Beyond quantitative improvements, DanceChat offers a new perspective on choreography modeling, where natural language serves not only as descriptive guidance but also as a bridge across modalities.